# Incorporating Preconditioning into Accelerated Approaches: Theoretical Guarantees and Practical Improvement

### v1

@@ 3.1 Objective function
Assume that $f(x)$ satisfies 2. Therefore, it holds that $\forall x\in\mathbb{R}^{d}$
@@ 3.2 Preconditioning
To prove Proposition 2, it is enough to apply the definition of the induced norm.
@@
Assume that the sequence of matrices $\{D_{k}\}_{k=-1}^{\infty}$ satisfies 3. If the update rule (2) has the form (5)/(6) with (7), then the following inequality holds:
@@ 4 Algorithms and Theoretical Results
Suppose that Assumptions 1, 2 and 3 hold. Then, after $K$ iterations of Algorithm 1 with $\gamma=\frac{(1-\beta_{1})^{2}e}{12L}$, we have $f\left(\frac{1}{W_{K-1}}\sum\limits_{k=0}^{K-1}w_{k}x_{k}\right)-f(x^{*})$ $\leq 4\exp\left(-\frac{(1-\beta_{1})\mu eK}{48L\Gamma}\right)L\left\lVert{x}_{0}-x^{*}\right\rVert_{\hat{D}_{-1}}^{2},$ where $W_{K-1}=\sum_{k=0}^{K-1}w_{k}$, $w_{k}=\left(1-\frac{\mu F}{4\Gamma}\right)^{-(k+1)}$.
@@
Under the conditions of Theorem 1, the required number of iterations $K$ of Algorithm 1 for reaching $\varepsilon$-accuracy, i.e. $f(x_{\text{out}})-f(x^{*})\leq\varepsilon$, can be upper bounded as $K=\mathcal{O}\left(\frac{L\Gamma}{\mu e(1-\beta_{1})}\log\left(\frac{L{\left\lVert x_{0}-x^{*}\right\rVert}^{2}}{\varepsilon}\right)\right).$
As for the Algorithm 2, the theorem of convergence can be formulated as follows.
@@
Suppose that Assumptions 1, 2 and 3 hold. Then, after $K$ iterations of Algorithm 2 with $\gamma_{k}\equiv\frac{e}{L},\xi_{k}\equiv\sqrt{\frac{L\Gamma}{\mu e}}$ and $\theta_{k}\equiv\frac{\sqrt{L\Gamma}}{\sqrt{\mu e}+\sqrt{L\Gamma}}$, we have $\left\lVert x^{k+1}-x^{*}\right\rVert_{\hat{D}_{k}}^{2}$ $\leq\exp\left(-K\sqrt{\frac{\mu e}{4L\Gamma}}\right)\left[\left\lVert x^{0}-x^{*}\right\rVert_{\hat{D}_{-1}}^{2}+\frac{2\Gamma}{\mu}(f(x^{0})-f(x^{*}))\right].$
@@
Under the conditions of Theorem 2, the required number of iterations $K$ of Algorithm 2 for reaching $\varepsilon$-accuracy, i.e. ${\left\lVert x^{k}-x^{*}\right\rVert}^{2}\leq\varepsilon$, can be upper bounded as $K=\mathcal{O}\left(\sqrt{\frac{L\Gamma}{\mu e}}\log\left(\frac{{\left\lVert x_{0}-x^{*}\right\rVert}^{2}+\frac{\Gamma}{\mu}(f(x^{0})-f(x^{*}))}{\varepsilon}\right)\right).$
@@ 5 Discussion of the results
    The induced norm. As a convergence criterion in Theorems 1 and 2, we employ the norm induced by the preconditioning matrix. Essentially, one can transition to the $L_{2}$-space by applying Proposition 2, which introduces a multiplicative factor $\Gamma$ related to the initial distance to the optimum. However, this factor is not critical, as it appears inside a logarithmic term in the iteration complexity bound on $K$.
@@
This phenomenon can be attributed to the fact that the scaling matrix allows one to operate over a modified landscape of the objective function, where, due to Proposition 2, the effective smoothness and strong convexity constants are altered accordingly; that is, $\mu$ becomes $\frac{\mu}{\Gamma}$, and $L$ becomes $\frac{L}{e}$.
@@ 6.1 Setup
Optimization methods. For our experiments, we implemented two optimization methods such as Heavy Ball, Nesterov and their scaled versions. Preconditioning matrix is chosen as for the Adam approach. For Algorithm 1, we chose a momentum parameter $\beta_{1}=0.9$, for Algorithm 2 we chose the hyperparameters according to the theory. Learning rate $\gamma$ is chosen as the best option after tuning.
@@ A.1 Proof of Preconditioned Heavy-Ball method
Suppose that Assumptions 1, 2 and 3 hold. Then, if the sequence $\{x_{k}\}_{k=0}$ is produced by Algorithm 1 and $\gamma$ satisfies $F:=\frac{\gamma}{1-\beta_{1}}\leq\frac{e}{4L},$ the next inequality holds: $\frac{F}{2}\left(f(x_{k})-f(x^{*})\right)$ $\leq\left(1-\frac{\mu F}{4\Gamma}\right)\left\lVert\tilde{x}_{k}-x^{*}\right\rVert_{\hat{D}_{k-1}}^{2}-\left\lVert\tilde{x}_{k+1}-x^{*}\right\rVert_{\hat{D}_{k}}^{2}$ $+\frac{3LF}{e}\left\lVert x_{k}-\tilde{x}_{k}\right\rVert_{\hat{D}_{k}}^{2}$ for all $k\geq 0$, where $\tilde{x}_{k}=x_{k}-\frac{\beta_{1}\gamma}{1-\beta_{1}}V_{k-1}$.
@@
To upper bound the second term in (A.1), we apply Proposition 1 and Proposition 2:
@@
For the third term in (A.1), according to 1, one can obtain
@@
For the last term in (A.1), we use the Young’s inequality (Proposition 4) with $\lambda=\frac{1}{2L}$ and Proposition 1:
@@
To bound the last term, we apply one-dimensional Young’s inequality (Proposition 4) with $\lambda=1$ to obtain
@@
where in the last inequality we applied $\frac{\mu}{\Gamma}\leq\frac{L}{e}$. Due to the bound on $F$ we have $\frac{2LF^{2}}{e}\leq\frac{F}{2}$. What is more, applying Proposition 3 with $\beta_{2}\geq 1-\frac{\mu F}{4\Gamma C}$ allows to get
@@
Suppose that Assumptions 2 and 3 hold. Thus, if the sequence $\{x_{k}\}_{k=0}$ is produced by Algorithm 1, the following inequality is satisfied for all $k$: $\left\lVert\tilde{x}_{k}-x_{k}\right\rVert_{\hat{D}_{k}}^{2}\leq\frac{2L\beta_{1}^{2}\gamma^{2}}{(1-\beta_{1})^{3}e}\sum\limits_{t=0}^{k-1}\beta_{1}^{k-1-t}\left(1+\frac{\mu F}{4\Gamma}\right)^{k-t}(f(x_{t})-f(x^{*})).$
@@
where we apply the Jensen’s inequality (Proposition 5) for the convex function $\parallel\cdot\parallel^{2}$. Moreover, due to Proposition 3 with $\beta_{2}\geq 1-\frac{\mu F}{4\Gamma C}$, we have
@@
where in the last inequality we applied Proposition 1 and Proposition 2. Substituting (A.1) into (A.1), we finish the proof. ∎
@@
Suppose that Assumptions 1, 2 and 3 hold. Then, after $K$ iterations of Algorithm 1 with $\gamma=\frac{(1-\beta_{1})^{2}e}{12L}$, we have $f\left(\frac{1}{W_{K-1}}\sum\limits_{k=0}^{K-1}w_{k}x_{k}\right)-f(x^{*})$ $\leq 4\exp\left(-\frac{(1-\beta_{1})\mu eK}{48L\Gamma}\right)L\left\lVert{x}_{0}-x^{*}\right\rVert_{\hat{D}_{-1}}^{2},$ where $W_{K-1}=\sum_{k=0}^{K-1}w_{k}$, $w_{k}=\left(1-\frac{\mu F}{4\Gamma}\right)^{-(k+1)}$.
@@
Let us denote $w_{k}=\left(1-\frac{\mu F}{4\Gamma}\right)^{-(k+1)}$. Therefore, summing the results from Lemma 1 with weights $w_{k}$, we get
@@
The main question arises for the last term. Applying Lemma 2, one can obtain
@@
Applying the Jensen’s inequality (Proposition 5) to the left-hand side and substituting the choice of $\gamma$, we conclude the proof. ∎
@@ A.2 Proof of Preconditioned Nesterov Method
Suppose that Assumptions 1, 2 and 3 hold. Then, if the sequence $\{x_{k}\}_{k=0}$ is produced by Algorithm 2, for all $k$ and any $u\in\mathbb{R}^{d}$ the following inequality is satisfied: $f(x_{f}^{k+1})$ $\leq f(u)-\left\langle\nabla f(x_{g}^{k}),u-x_{g}^{k}\right\rangle-\frac{\mu}{2}\left\lVert u-x_{g}^{k}\right\rVert_{2}^{2}$ $+\left(\frac{L\gamma_{k}}{2e}-1\right)\gamma_{k}\left\lVert\nabla f(x_{g}^{k})\right\rVert^{2}_{\hat{D}_{k}^{-1}}.$
@@
Applying Proposition 2 to the last term, we obtain
@@
The first term can be upper bounded with the 1:
@@
Suppose that Assumptions 1, 2 and 3 hold. Then, if the sequence $\{x_{k}\}_{k=0}$ is produced by Algorithm 2, for all $k$ with $\xi_{k}\geq 1,\frac{\xi_{k}^{2}\gamma_{k}\mu}{\Gamma}\geq 1$, $\gamma_{k}\leq\frac{e}{L}$ and $\theta_{k}=\frac{\xi_{k}}{1+\xi_{k}}$, the next inequality is satisfied: $\left\lVert x^{k+1}-x^{*}\right\rVert_{\hat{D}_{k}}^{2}$ $+2\gamma_{k}\xi_{k}^{2}(f(x_{f}^{k+1})-f(x^{*}))$ $\leq\left(1-\frac{1}{\xi_{k}}\right)\left[\left\lVert x^{k}-x^{*}\right\rVert_{\hat{D}_{k}}^{2}+2\gamma_{k}\xi_{k}^{2}(f(x_{f}^{k})-f(x^{*}))\right].$
@@
From Lemma 3 with $u=x_{f}^{k}$ and $u=x^{*}$, under Proposition 2 we get
@@
Suppose that Assumptions 1, 2 and 3 hold. Then, after $K$ iterations of Algorithm 2 with $\gamma_{k}\equiv\frac{e}{L},\xi_{k}\equiv\sqrt{\frac{L\Gamma}{\mu e}}$ and $\theta_{k}\equiv\frac{\sqrt{L\Gamma}}{\sqrt{\mu e}+\sqrt{L\Gamma}}$, we have $\left\lVert x^{k+1}-x^{*}\right\rVert_{\hat{D}_{k}}^{2}$ $\leq\exp\left(-K\sqrt{\frac{\mu e}{4L\Gamma}}\right)\left[\left\lVert x^{0}-x^{*}\right\rVert_{\hat{D}_{-1}}^{2}+\frac{2\Gamma}{\mu}(f(x^{0})-f(x^{*}))\right].$
@@
Starting with Lemma 4:
@@
We use the Proposition 3 with $\beta_{2}\geq 1-\frac{1}{2C}\sqrt{\frac{\mu e}{L\Gamma}}$ such that
@@
For Theorem 2 it is enough to apply Proposition 2 to obtain the convergence in $L_{2}$-space.

### v2

@@ 3.1 Objective function
Assume that $f(x)$ satisfies ˜2. Therefore, it holds that $\forall x\in\mathbb{R}^{d}$
@@ 3.2 Preconditioning
To prove Proposition˜2, it is enough to apply the definition of the induced norm.
@@
Assume that the sequence of matrices $\{D_{k}\}_{k=-1}^{\infty}$ satisfies ˜3. If the update rule (2) has the form (5)/(6) with (7), then the following inequality holds:
@@ 4 Algorithms and Theoretical Results
Suppose that Assumptions 1, 2 and 3 hold. Then, after $K$ iterations of Algorithm˜1 with $\gamma=\frac{(1-\beta_{1})^{2}e}{12L}$, we have $f\left(\frac{1}{W_{K-1}}\sum\limits_{k=0}^{K-1}w_{k}x_{k}\right)-f(x^{*})$ $\leq 4\exp\left(-\frac{(1-\beta_{1})\mu eK}{48L\Gamma}\right)L\left\lVert{x}_{0}-x^{*}\right\rVert_{\hat{D}_{-1}}^{2},$ where $W_{K-1}=\sum_{k=0}^{K-1}w_{k}$, $w_{k}=\left(1-\frac{\mu F}{4\Gamma}\right)^{-(k+1)}$.
@@
Under the conditions of Theorem˜1, the required number of iterations $K$ of Algorithm˜1 for reaching $\varepsilon$-accuracy, i.e. $f(x_{\text{out}})-f(x^{*})\leq\varepsilon$, can be upper bounded as $K=\mathcal{O}\left(\frac{L\Gamma}{\mu e(1-\beta_{1})}\log\left(\frac{L{\left\lVert x_{0}-x^{*}\right\rVert}^{2}}{\varepsilon}\right)\right).$
As for the Algorithm˜2, the theorem of convergence can be formulated as follows.
@@
Suppose that Assumptions 1, 2 and 3 hold. Then, after $K$ iterations of Algorithm˜2 with $\gamma_{k}\equiv\frac{e}{L},\xi_{k}\equiv\sqrt{\frac{L\Gamma}{\mu e}}$ and $\theta_{k}\equiv\frac{\sqrt{L\Gamma}}{\sqrt{\mu e}+\sqrt{L\Gamma}}$, we have $\left\lVert x^{k+1}-x^{*}\right\rVert_{\hat{D}_{k}}^{2}$ $\leq\exp\left(-K\sqrt{\frac{\mu e}{4L\Gamma}}\right)\left[\left\lVert x^{0}-x^{*}\right\rVert_{\hat{D}_{-1}}^{2}+\frac{2\Gamma}{\mu}(f(x^{0})-f(x^{*}))\right].$
@@
Under the conditions of Theorem˜2, the required number of iterations $K$ of Algorithm˜2 for reaching $\varepsilon$-accuracy, i.e. ${\left\lVert x^{k}-x^{*}\right\rVert}^{2}\leq\varepsilon$, can be upper bounded as $K=\mathcal{O}\left(\sqrt{\frac{L\Gamma}{\mu e}}\log\left(\frac{{\left\lVert x_{0}-x^{*}\right\rVert}^{2}+\frac{\Gamma}{\mu}(f(x^{0})-f(x^{*}))}{\varepsilon}\right)\right).$
@@ 5 Discussion of the results
    The induced norm. As a convergence criterion in Theorems 1 and 2, we employ the norm induced by the preconditioning matrix. Essentially, one can transition to the $L_{2}$-space by applying Proposition˜2, which introduces a multiplicative factor $\Gamma$ related to the initial distance to the optimum. However, this factor is not critical, as it appears inside a logarithmic term in the iteration complexity bound on $K$.
@@
This phenomenon can be attributed to the fact that the scaling matrix allows one to operate over a modified landscape of the objective function, where, due to Proposition˜2, the effective smoothness and strong convexity constants are altered accordingly; that is, $\mu$ becomes $\frac{\mu}{\Gamma}$, and $L$ becomes $\frac{L}{e}$.
@@ 6.1 Setup
Optimization methods. For our experiments, we implemented two optimization methods such as Heavy Ball, Nesterov and their scaled versions. Preconditioning matrix is chosen as for the Adam approach. For Algorithm˜1, we chose a momentum parameter $\beta_{1}=0.9$, for Algorithm˜2 we chose the hyperparameters according to the theory. Learning rate $\gamma$ is chosen as the best option after tuning.
@@ A.1 Proof of Preconditioned Heavy-Ball method
Suppose that Assumptions 1, 2 and 3 hold. Then, if the sequence $\{x_{k}\}_{k=0}$ is produced by Algorithm˜1 and $\gamma$ satisfies $F:=\frac{\gamma}{1-\beta_{1}}\leq\frac{e}{4L},$ the next inequality holds: $\frac{F}{2}\left(f(x_{k})-f(x^{*})\right)$ $\leq\left(1-\frac{\mu F}{4\Gamma}\right)\left\lVert\tilde{x}_{k}-x^{*}\right\rVert_{\hat{D}_{k-1}}^{2}-\left\lVert\tilde{x}_{k+1}-x^{*}\right\rVert_{\hat{D}_{k}}^{2}$ $+\frac{3LF}{e}\left\lVert x_{k}-\tilde{x}_{k}\right\rVert_{\hat{D}_{k}}^{2}$ for all $k\geq 0$, where $\tilde{x}_{k}=x_{k}-\frac{\beta_{1}\gamma}{1-\beta_{1}}V_{k-1}$.
@@
To upper bound the second term in (A.1), we apply Proposition˜1 and Proposition˜2:
@@
For the third term in (A.1), according to ˜1, one can obtain
@@
For the last term in (A.1), we use the Young’s inequality (Proposition˜4) with $\lambda=\frac{1}{2L}$ and Proposition˜1:
@@
To bound the last term, we apply one-dimensional Young’s inequality (Proposition˜4) with $\lambda=1$ to obtain
@@
where in the last inequality we applied $\frac{\mu}{\Gamma}\leq\frac{L}{e}$. Due to the bound on $F$ we have $\frac{2LF^{2}}{e}\leq\frac{F}{2}$. What is more, applying Proposition˜3 with $\beta_{2}\geq 1-\frac{\mu F}{4\Gamma C}$ allows to get
@@
Suppose that Assumptions 2 and 3 hold. Thus, if the sequence $\{x_{k}\}_{k=0}$ is produced by Algorithm˜1, the following inequality is satisfied for all $k$: $\left\lVert\tilde{x}_{k}-x_{k}\right\rVert_{\hat{D}_{k}}^{2}\leq\frac{2L\beta_{1}^{2}\gamma^{2}}{(1-\beta_{1})^{3}e}\sum\limits_{t=0}^{k-1}\beta_{1}^{k-1-t}\left(1+\frac{\mu F}{4\Gamma}\right)^{k-t}(f(x_{t})-f(x^{*})).$
@@
where we apply the Jensen’s inequality (Proposition˜5) for the convex function $\parallel\cdot\parallel^{2}$. Moreover, due to Proposition˜3 with $\beta_{2}\geq 1-\frac{\mu F}{4\Gamma C}$, we have
@@
where in the last inequality we applied Proposition˜1 and Proposition˜2. Substituting (A.1) into (A.1), we finish the proof. ∎
@@
Suppose that Assumptions 1, 2 and 3 hold. Then, after $K$ iterations of Algorithm˜1 with $\gamma=\frac{(1-\beta_{1})^{2}e}{12L}$, we have $f\left(\frac{1}{W_{K-1}}\sum\limits_{k=0}^{K-1}w_{k}x_{k}\right)-f(x^{*})$ $\leq 4\exp\left(-\frac{(1-\beta_{1})\mu eK}{48L\Gamma}\right)L\left\lVert{x}_{0}-x^{*}\right\rVert_{\hat{D}_{-1}}^{2},$ where $W_{K-1}=\sum_{k=0}^{K-1}w_{k}$, $w_{k}=\left(1-\frac{\mu F}{4\Gamma}\right)^{-(k+1)}$.
@@
Let us denote $w_{k}=\left(1-\frac{\mu F}{4\Gamma}\right)^{-(k+1)}$. Therefore, summing the results from Lemma˜1 with weights $w_{k}$, we get
@@
The main question arises for the last term. Applying Lemma˜2, one can obtain
@@
Applying the Jensen’s inequality (Proposition˜5) to the left-hand side and substituting the choice of $\gamma$, we conclude the proof. ∎
@@ A.2 Proof of Preconditioned Nesterov Method
Suppose that Assumptions 1, 2 and 3 hold. Then, if the sequence $\{x_{k}\}_{k=0}$ is produced by Algorithm˜2, for all $k$ and any $u\in\mathbb{R}^{d}$ the following inequality is satisfied: $f(x_{f}^{k+1})$ $\leq f(u)-\left\langle\nabla f(x_{g}^{k}),u-x_{g}^{k}\right\rangle-\frac{\mu}{2}\left\lVert u-x_{g}^{k}\right\rVert_{2}^{2}$ $+\left(\frac{L\gamma_{k}}{2e}-1\right)\gamma_{k}\left\lVert\nabla f(x_{g}^{k})\right\rVert^{2}_{\hat{D}_{k}^{-1}}.$
@@
Applying Proposition˜2 to the last term, we obtain
@@
The first term can be upper bounded with the ˜1:
@@
Suppose that Assumptions 1, 2 and 3 hold. Then, if the sequence $\{x_{k}\}_{k=0}$ is produced by Algorithm˜2, for all $k$ with $\xi_{k}\geq 1,\frac{\xi_{k}^{2}\gamma_{k}\mu}{\Gamma}\geq 1$, $\gamma_{k}\leq\frac{e}{L}$ and $\theta_{k}=\frac{\xi_{k}}{1+\xi_{k}}$, the next inequality is satisfied: $\left\lVert x^{k+1}-x^{*}\right\rVert_{\hat{D}_{k}}^{2}$ $+2\gamma_{k}\xi_{k}^{2}(f(x_{f}^{k+1})-f(x^{*}))$ $\leq\left(1-\frac{1}{\xi_{k}}\right)\left[\left\lVert x^{k}-x^{*}\right\rVert_{\hat{D}_{k}}^{2}+2\gamma_{k}\xi_{k}^{2}(f(x_{f}^{k})-f(x^{*}))\right].$
@@
From Lemma˜3 with $u=x_{f}^{k}$ and $u=x^{*}$, under Proposition˜2 we get
@@
Suppose that Assumptions 1, 2 and 3 hold. Then, after $K$ iterations of Algorithm˜2 with $\gamma_{k}\equiv\frac{e}{L},\xi_{k}\equiv\sqrt{\frac{L\Gamma}{\mu e}}$ and $\theta_{k}\equiv\frac{\sqrt{L\Gamma}}{\sqrt{\mu e}+\sqrt{L\Gamma}}$, we have $\left\lVert x^{k+1}-x^{*}\right\rVert_{\hat{D}_{k}}^{2}$ $\leq\exp\left(-K\sqrt{\frac{\mu e}{4L\Gamma}}\right)\left[\left\lVert x^{0}-x^{*}\right\rVert_{\hat{D}_{-1}}^{2}+\frac{2\Gamma}{\mu}(f(x^{0})-f(x^{*}))\right].$
@@
Starting with Lemma˜4:
@@
We use the Proposition˜3 with $\beta_{2}\geq 1-\frac{1}{2C}\sqrt{\frac{\mu e}{L\Gamma}}$ such that
@@
For Theorem 2 it is enough to apply Proposition˜2 to obtain the convergence in $L_{2}$-space.